On 31 August, M traders purchased a machine on credit for a cost price of R402 500 (including VAT at 15%). Machine is depreciated over 6 years according to the straight-line method and there is residual value on this machine. M traders has a 30 June year end. Calculate depreciation and accumulated depreciation for 2019 and 2020
1
Expert's answer
2020-10-13T10:45:13-0400

Cost of machine without VAT:

"\frac{402,500}{1.15}=350,000"

The depreciation per month:

"\frac{350,000}{6*12}=4861.11"

The accumulated depreciation for 2019:

"4861.11*10=48,611.1"

The accumulated depreciation for 2020:

"4861.11*22=106,944.42"